OpenStudy (anonymous):

{4x+2y=4 {6x+2y=8 And Equation by Elimination

OpenStudy (anonymous):

an*

OpenStudy (anonymous):

Try subtracting the first equation from the second... Eq # 1: 4x+2y=4 Eq # 2: 6x+2y=8 Eq # 2 - Eq # 1 (Eq2) 6x + 2y = 8 (Eq1) - (4x + 2y = 8) ---------------------

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Hope it makes sense now :) glad to help!